Journalist Report – October 28th

Monday 28th October Sol 1

After a freezing night (literally) we awake to a cold but clear day. The mountain to the south (Henry Mountains) is wreathed in cloud, and shows a glazing of snow. Today is our first day ‘on Mars’! This takes a bit of psychological adjustment.

We congratulate each other on a successful ‘journey’, and get on with the day. This is mainly about bedding in systems: getting the radios working, making equipment for the micrometeorite experiment, setting up the fire blanket for the kitchen.

The radios were a bit of a mess, with a number not working in different ways, but we were able to get a working minimum set functioning eventually.

Today we are expecting our 5th team member, David, to arrive, or ‘descend in the latest shuttle’. David is very experienced in the MDRS, so will be able to train us up in the use of the suits, etc.

The afternoon shift was spent working on equipment for the micrometeorite expedition, tying and marking string in 1 metre lengths for the grid to be set up later in the field.
After another delicious re-constituted meal, we move onto the CapCom reports and wind up the day.

Operations Report – October 28th

Name of person filing report: David Mateus
Non-nominal systems: Opportunity
Notes on non-nominal systems: Its brakes do not work
Generator: run
Hours run: 8.5
From what time last night: 18:00
To what time this morning: 2:30
List any additional daytime hours when the generator was run: none
Solar— SOC % (Before generator is run at night): 84%
Diesel Reading –75 %
Station Propane Reading – 88 %
Ethanol Free Gasoline – N/A
Water (loft tank): 47 gallons
Water Meter: 1454197 units
Water (static tank) – 350 gallons
Static to Loft Pump used – Yes
Water in GreenHab – 0 gallons
Water in ScienceDome: 0 gallons
Toilet tank emptied: yes

Deimos rover used: It is not in the hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: No
Hours: (before EVA): 105.5 h
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging):100%
Currently charging: Yes

Opportunity rover used: No
Hours: 65.5 h
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 111.5
Beginning charge: 100%
Ending charge:100%
Currently charging: Yes

Notes on rovers: No

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) No
Reason for use: None
Oil Added? No
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? Used by assistant directors to go to town

CrewCar used and why, where? Used to go to Grand Junction

General notes and comments: None
Summary of internet: functional
Summary of suits and radios: No used
Summary of Hab operations: All systems are nominal
Summary of GreenHab operations: No used
Summary of ScienceDome operations: All systems are nominal, used for experiment preparations
Summary of RAM operations: No used
Summary of any observatory issues: Re-centered the focuser by the assistant directors
Summary of health and safety issues: All the crew are healthy
Questions, concerns and requests to Mission Support: We need new batteries for six radios.

Sol Summary – October 28th

Sol:1

Summary Title: Touchdown on Mars

Author’s name: Guy Murphy

Mission Status: The crew woke up on Mars this morning, this being our first day in full simulation mode.

Sol Activity Summary: The morning was dedicated to preparing and commencing use of the radios, as well as other minor maintenance, cleaning and administrative tasks. In the afternoon our 5th crew member David Mateus arrived. Crew geologist Andrew Wheeler continued preparation of equipment for the micro-meteorite collection study. Three cooked meals were prepared today.

Look Ahead Plan: Our goal is to commence our EVA schedule tomorrow, with David Mateus leading us on our first EVAs to learn about the procedures and equipment.

Anomalies in work: Not all the radios were found to be working.

Weather: Very cold overnight, clear and sunny daytime sky.

Crew Physical Status: All crew in good health.

EVA: Not commenced today.

Reports to be filed: Sol Summary, Operations Report, EVA Requests (2), Journalists Report, Science Report, Days Photos, Crew Car Collection Report, Mission Patch.

Support Requested: New batteries were requested to supplement the existing limited supply of working batteries for future crews.

Journalist Report Oct 27th

Journalist Report October 27

Journalist: Sandy Dance

After travelling the world the team members finally assemble and meet each other, mostly for the first time, at the Days Inn in Grand Junction Colorado. We spend an afternoon buying all those last minute things we require for 2 or 4 weeks in the wilderness. And incidentally get to know each other.

We are a good team: Diane the nutritionist, Andrew the geologist, Sandy the IT and generalist, and our Commander Guy, the heritage consultant. Unfortunately the 5th person we were expecting had a medical withdrawal and couldn’t make it. Luckily Shannon was able to donate one of her interns to the team, David who studies mechatronics at Peru University.

That afternoon we drive the 2 ½ hours to Hanksville. There we meet up with Shannon in the hot and dusty little, but strangely attractive town. This is at the Rock Shop run by one Cathy, a funky little place with a vast array of geological specimens, especially dinosaur coprolites!

Finally we drive to the MRDS down a dusty little track winding around barren hillocks and crazy shaped rocks, turn a corner, and there it is: the Mars habitat.

We spend the rest of the time shifting supplies into the hab and tidying things up.

After a knockup meal, we collapse into bed, each with our tiny but lovable “stateroom”.

Sunday 27th October

We wake at 7am for a hearty breakfast of coffee, tea and porridge, then knuckle down to a morning of cleaning the hab and base, charging and preparing the spacesuits, checking the radios, rearranging the furniture, and moving rubbish from around the hab out of site.

The weather turns cold, cloudy and windy, not like yesterday.

Guy and I manage to skive off at one point and take a walk up the hill, something we can do today as we are not yet in ‘sim’. The walk up is over a hard crust of clay lying on top of very soft and dry clay powder. Our footsteps break through if we’re not careful, we don’t want to leave a permanent trace. The surface is rendered by rain and drought into a sort of elephants hide of minute erosion texture. We make it to the top and are greeted with a magnificent panorama around the horizon: eroded hills, buttes, mountains in the distance, all in a palette of pinks, greys and greens. One feature, Factory Butte, looks like its name, but with a skirt of eroded scree at 45 degrees.

Another knockup meal, lunch prepared from the dry goods (we are going to get used to this): cheesy dumplings with thick soup, just the thing!

Science Report Oct 27th

Crew 214 Science Report 27-Oct-2019

Submitted by Science Officer Andrew Wheeler

1. EVA Suit Maintenance: The hab has been set up following arrival. All backpacks have been plugged in and fully charged. Helmets are clean and contained in their cloth bags except for one helmet, for which, the cloth bag is missing. Radios have been tested and are fully charged. Communication earpieces remain to be tested for faulty performance.

2. Green Hab: The green hab has been activated and is awaiting water and additional soil base before sowing of this seasons crops.

3. Astronomy Dome Operations: No astronomical observations have been planned for this crew rotation.

4. Science Dome Operations: The Science Dome has been activated and is awaiting EVAs for sample acquisition and sample testing.

5. RAM Operations: The RAM Dome has been activated and is ready for contingencies.

6. EVA: No preparations or activities have been undertaken to date. The access routes have been identified to allow sample locations to be identified.

7. Closed Loop Food Waste Study: This study has been initiated and is ongoing. Day’s waste: 79g; consisting of 323 calories, 10g fat, 5g saturated fat, 1179mg sodium, 52g carbohydrates, 7g fibre, 6g sugar, 7g protein.

8. Additional activities: Nil

Sol Summary Report Oct 27th

Sol:0

Summary Title: Preparing for Landing

Author’s name: Guy Murphy

Mission Status: Since its arrival yesterday, Crew 214 of Expedition Boomerang has been adjusting to its new surroundings and preparing the campus for commencement of full simulation mode.

Sol Activity Summary: Today we cleaned the bottom level of the habitat, and the outside of the buildings. We continued to familiarise ourselves with the various operational systems and procedures. We prepared 3 cooked meals.

Look Ahead Plan: We will commence full simulation mode at first light tomorrow morning. Our fifth crew member David will join us early afternoon. David will brief the crew about using the suits and EVA procedure.

Anomalies in work: N/A

Weather: Clear skies all day, cooler weather later in the afternoon.

Crew Physical Status: All crew in good health, adjusting to MDRS conditions.

EVA: EVA schedule not commenced.

Reports to be filed: Sol Summary, Science Report, Journalists Report, HSO Pre-Mission Checklist, Photo of the day

Support Requested: We welcome feedback from Mission Support as to how we can improve our practise of comms procedures.

Journalist Report – October 26th

Journalist’s Report

Saturday 26th October

After travelling the world the team members finally assemble and meet each other, mostly for the first time, at the Days Inn in Grand Junction Colorado. We spend an afternoon buying all those last minute things we require for 2 or 4 weeks in the wilderness. And incidentally get to know each other.

We are a good team: Dianne the nutritionist, Andrew the geologist, Sandy the IT and generalist, and our Commander Guy, the heritage consultant. Unfortunately the 5th person we were expecting had a medical withdrawal and couldn’t make it. Luckily Shannon was able to donate one of her interns to the team, David who studies mechatronics at the National University of Colombia.

That afternoon we drive the 2 ½ hours to Hanksville. There we meet up with Shannon in the hot and dusty little, but strangely attractive town. This is at the Rock Shop run by one Cathy, a funky little place with a vast array of geological specimens, especially dinosaur coprolites!

Finally we drive to the MDRS down a dusty little track winding around barren hillocks and crazy shaped rocks, turn a corner, and there it is: the Mars habitat.

We spend the rest of the time shifting supplies into the Hab and tidying things up.

After a knock-up meal, we collapse into bed, each with our tiny but lovable “stateroom”.